OpenStudy (anonymous):

Y varies inversely with x. If y = 21 when x = 22, find x when y = 23

OpenStudy (johnweldon1993):

y = k/x we have 'y' and we have 'x' so we can solve for 'k' 21 = k/22 22 * 21 = 462 Now that we have 'k' we can solve for x in the other equation y = k/x 23 = 462/x x = 462 / 23 x = ...

OpenStudy (anonymous):

If y varies with x inversely => that x*y is a constant quantity ( That means x1*y1=x2*y2=x3*y3 ... and so-on) here x1=22,y1=21y2=23 then u have to find x2 use x1*y1=x2*y2 Put the values u will get the Ans

OpenStudy (anonymous):

is the answer 20.1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you!